### v2

# Enhancing Dynamic CT Image Reconstruction with Neural Fields and Optical Flow

Abstract
In this paper, we investigate image reconstruction for dynamic Computed Tomography. The motion of the target with respect to the measurement acquisition rate leads to highly resolved in time but highly undersampled in space measurements. Such problems pose a major challenge: not accounting for the dynamics of the process leads to a poor reconstruction with non-realistic motion. Variational approaches that penalize time evolution have been proposed to relate subsequent frames and improve image quality based on classical grid-based discretizations. Neural fields have emerged as a novel way to parameterize the quantity of interest using a neural network with a low-dimensional input, benefiting from being lightweight, continuous, and biased towards smooth representations. The latter property has been exploited when solving dynamic inverse problems with neural fields by minimizing a data-fidelity term only. We investigate and show the benefits of introducing explicit motion regularizers for dynamic inverse problems based on partial differential equations, namely, the optical flow equation, for the optimization of neural fields. We compare it against its unregularized counterpart and show the improvements in the reconstruction. We also compare neural fields against a grid-based solver and show that the former outperforms the latter in terms of PSNR in this task.
keywords:
Dynamic Computed Tomography, Neural fields, Physics-Informed Neural Networks, Optical flow
1 Introduction
In many imaging tasks, the target object changes during the data acquisition. In clinical settings for instance, imaging techniques such as Computed Tomography (CT), Positron Emission Tomography (PET) or Magnetic Resonance Imaging (MRI) are used to study moving organs such as the heart or the lungs. Usually, the acquired data is a time series collected at several finely discretized times $0=t_{1}<\ldots<t_{N_{T}}=T$. However, the motion of these organs prevents the scanners from taking enough measurements at a single time instance, resulting in highly undersampled spatial measurements. A naive way to proceed is by neglecting the time component and solving several static inverse problems. However, the lack of information makes this frame-by-frame reconstruction a severely ill-posed problem leading to a poor reconstruction. A common procedure is to bin the data in time, where several time-step measurements are collapsed into one to gain more information in space at the cost of losing temporal resolution and introducing artefacts in the reconstruction. It is therefore necessary to seek a spatiotemporal quantity with coherence between subsequent frames whose reconstruction considers the dynamics of the process.
1.1 Dynamic Inverse Problems
In (static) inverse problems, we aim to reconstruct a quantity $u:\Omega\subset\mathbb{R}^{d}\to\mathbb{R}$ from a discrete set of measurements $f$ obtained by some device by solving an equation of the form
$Ku+\varepsilon=f.$ (1)
Here $K$ is the forward operator that models the imaging process by mapping the continuous object $u$ into a discrete set of measurements, and $\varepsilon$ is the noise coming from the measurement acquisition.
For dynamic inverse problems, we formulate the problem as follows: we let the measurements $(f_{t})_{t\in\{t_{1},\ldots,t_{N_{T}}\}}\subset\mathbb{R}^{M}$, with $M$ the number of measurements collected at a given time. The goal is to recover a time-dependent quantity $u:\Omega_{T}:=\Omega\times[0,T]\to\mathbb{R}$ by solving the equation below
$K_{t}[u_{t}]+\varepsilon_{t}=f_{t},\quad\text{ for }t\in\{t_{1},\ldots,t_{N_{T}}\}.$ (2)
Here $u_{t}$, $K_{t}[u_{t}]\in\mathbb{R}^{M}$, and $\varepsilon_{t}$ are the solution, the imaging process, and the noise at time $t$, respectively.
The classical way to address this problem is to discretize the solution using a grid-based representation $u\in\mathbb{R}^{N\times N_{T}}$, the Casorati matrix, with $N$ the number of pixels in space and $N_{T}$ the number of frames. The columns of this matrix represent the solution at the corresponding time step. The problem is then solved using a variational formulation which consists of a data-fidelity term plus some suitable regularizer $\mathcal{R}$ that seeks correlation between the columns of $u$. Common examples of such regularizers include sparsity-based regularizers, inspired by the idea that the sought quantity can be compactly represented on a suitable basis, e.g., total variation, wavelets or shearlets, see [1, 2, 3, 4]; the nuclear norm, which promotes the solution to be a low-rank matrix, see [5]; first-order time derivative penalizers, suitable for sequences with small displacements, see [6, 7].
Another option is explicitly considering the target’s motion by including its velocity field $\mathbb{v}$ in the formulation. We call this a motion-based regularizer, where a partial differential equation (PDE) $r(u,\mathbb{v})=0$, relating $u$, the velocity field $\mathbb{v}$, and their derivatives, is used to impose a physical prior. Usual examples include the optical flow and the continuity equations. In this paper, we focus on the former:
$r(u,\mathbb{v}):=\partial_{t}u+\mathbb{v}\cdot\nabla u=0,\quad\text{ in }\Omega_{T}.$ (3)
This equation shall be discussed in more detail in section 2.1. As the underlying motion is unknown the overall problem is commonly referred to as a joint image reconstruction and motion estimation task, see [8, 9, 10, 11]. We refer to [12] for an extensive review of dynamic inverse problems.
Classical grid-based representations of the spatiotemporal image suffer from two issues: (1) their lack of regularity which motivates the use of several regularizers such as the ones mentioned above, and (2) their complexity grows exponentially with the dimension and polynomially with the discretization due to the curse of dimensionality which can incur in memory burden. The latter is particularly relevant to large-scale problems, for instance, when the measurement frame rate and/or spatial resolution is high, in 3D+time domains, etc. In the next section, we introduce neural fields, an alternative continuous representation using deep neural networks.
1.2 Neural Fields
In recent years, coordinate-based multilayer perceptrons (MLPs) have been employed as a new way of parameterizing quantities of interest. In computer vision, these are referred to as neural fields or implicit neural representations [13, 14], while the term Physics-Informed Neural Networks (PINNs) has been adopted when used to solve PDEs [15, 16]. The main idea is to use a neural network $u_{\theta}$ with trainable weights $\theta$ as an ansatz for the solution of the problem. It takes as input a spatio-temporal point $(x,t)\in\Omega_{T}$, and outputs the value $u_{\theta}(x,t)$ at that point, e.g., the intensity of the image at that particular time and location. The problem is then rephrased as a non-convex optimization that seeks optimal weights $\theta$. The method requires training a neural network for every new instance, thus, it is said to be self-supervised and differs from the usual learning framework where a solution map is found by training a network over large datasets. Applications of neural fields include image reconstruction in CT [17, 18, 19], MRI [20, 21, 22, 23, 24], image registration [25, 26, 27], continuous shape representation via signed distance functions [28], view synthesis [29], among others.
It is well-known that, under mild conditions, neural networks can approximate functions at any desired tolerance [30], but other properties have justified their widespread use:
  1. 1.
    Implicit regularization. Numerical experiments and theoretical results show that neural fields tend to learn smooth functions early during training, commonly referred to as spectral bias [31, 32, 33]. This is both advantageous and disadvantageous: neural fields can capture smooth regions of natural images but will struggle at capturing sharp edges. The latter can be overcome with Fourier feature encoding [34] or with sinusoidal activation functions as in SIREN [14]. We highlight that smoothness in time is highly desirable for dynamic inverse problems.
  2. 2.
    Overcoming the curse of dimensionality. In [35, 36] it is shown that the amount of weights needed to approximate the solution of particular PDEs grows polynomially on the dimension of the domain. For the same reason, only a few weights can represent complex images, leading to a lightweight, compact and memory-efficient representation. This has been exploited, for instance, in [37] for image compression.
  3. 3.
    Continuous and differentiable representation. This property is exploited in PINNs where the derivatives of a PDE are computed using automatic differentiation with an accuracy only limited by machine precision. In particular, the parametrization does not rely on a mesh as in finite differences or finite elements.
1.3 Contributions
In the context of dynamic inverse problems and neural fields, most of the literature relies entirely on the smoothness introduced by the network on the spatial and temporal variables to get a regularized solution. This motivates the minimization of the data-fidelity term without considering any explicit regularizers. Examples can be found on dynamic cardiac MRI in [21, 24], where the network outputs the real and imaginary parts of the signal, while in [22] the neural field is used to fit the measurements directly and then inference is performed by rendering the neural field at a uniform grid in the k-space followed by the inverse Fourier transform. In [38, 39] neural fields are used to solve a photoacoustic tomography dynamic reconstruction emphasizing their memory efficiency. Dynamic CT has been addressed in 3D+time domains in [19, 40], where the neural field parametrizes the initial frame and a deformation vector field warps it to get the subsequent frames in time. To the best of our knowledge, the previous are the only works using neural fields and a motion model via a deformable template.
Motivated by [8, 10] and leveraging the arbitrary resolution and automatic differentiation of neural fields to compute spatial and temporal derivatives, in this paper, we study neural fields in the context of dynamic inverse problems in a highly undersampled measurement regime with the optical flow equation as an explicit PDE-based motion regularizer imposed as a soft constraint as in PINNs. We do not consider the nuclear norm and the sparsity-based regularizers previously mentioned since they act on a discrete representation of the solution on a cartesian grid and hence do not exploit the mesh-free nature of the neural field: to use them, the neural field would need to be queried at points on a cartesian grid to get a discrete representation over which the regularizer can act. This has been done, for instance, in [23] for dynamic MRI.
Our findings are based on numerical experiments on dynamic CT performed on two synthetic and one real datasets, all in a 2D+time domain. The contributions are summarized as follows:
  • Constraining the neural field with an explicit motion model ensures that only physically feasible solution manifolds are considered. We demonstrate that our approach improves the reconstruction when compared to a motionless unregularized model.
  • We show how to leverage the mesh-free nature of neural fields to impose regularizers for imaging tasks.
  • We study the reconstruction obtained by neural fields and by a grid-based representation. We demonstrate that neural fields can outperform classical discretizations in terms of the quality of the reconstruction for highly undersampled dynamic CT.
The paper is organized as follows: in section 2 we introduce dynamic compute tomography, motion models with the optical flow equation, and the joint image reconstruction and motion estimation variational problem as in [8]; in section 3 we state the main variational problem to be minimized and study how to solve it with neural fields and with a grid-based representation; in section 4 we describe the datasets; in section 5 we investigate our method numerically and show its improvements in comparison to unregularized neural fields and the grid-based representation; we finish with the conclusions in section 6.
2 Dynamic Inverse Problems for Computed Tomography
In X-ray CT [41, 42, 43] the unknown $u$ is the non-negative absorption of photons of the imaged object. The forward model is a line integral whose precise formulation depends on the scanner: common geometries include parallel and fan beams. We focus on the latter, where the X-rays are emitted from several source points $p\in\{p_{1},\ldots,p_{P}\}\subset\mathbb{R}^{d}$ through the object $O$ in different directions towards $M$ sensors that collect the photons, see figure 1. Given a source point $p\in\mathbb{R}^{d}$, a set $\{L^{p}_{1},\ldots,L^{p}_{M}\}$ of $M$ lines is constructed with $L^{p}_{j}$ going from $p$ to the $j$-th sensor. The projection from $p$ can be represented as follows:
[…]
leading to $f\in\mathbb{R}^{M\times P}$ measurements when considering all source positions $p_{1},\ldots,p_{P}$.
For dynamic CT we consider both the object’s motion and the scanner’s rotation around it. Hence, the forward operator $K_{t}$ in (2) is the projection from a source point $p_{t}$:
$K_{t}[u_{t}]:=K[u_{t}](p_{t}).$
When the motion of $u$ is slow compared to the rotation of the scanner, it is possible to pose equation (2) as (1) by neglecting the time variable. In this case, $u$ can be reconstructed from the binned vector $f_{\text{bin}}:=[f_{t_{1}},\ldots,f_{t_{N_{T}}}]\in\mathbb{R}^{M\times N_{T}}$. To highlight the necessity of motion models, two naive reconstructions obtained using filtered back projection are shown in figure 2. The first row depicts the two-square phantom we will introduce in section 4 for our numerical experiments, and the measurements. The reconstructions are shown at the bottom. As expected, we cannot get a reliable reconstruction from one projection only. The image on the bottom right corresponds to the reconstruction from $f_{\text{bin}}$. The result is an image that blurs the motion of the squares.
In the next section, we introduce motion models to regularize the dynamic inverse problem.
[…]
Figure 1: Fan-beam geometry for CT. X-rays emitted from the source point $p$ in different directions are attenuated by the object and then measured in the sensors represented by the red dots.
Figure 2: Top row: ground truth image $u$ at frames 35 and 55 (out of 100) with two squares moving, and measurements $f\in\mathbb{R}^{M\times N_{T}}$. The red boxes indicate the projection obtained in frames 35 and 55. Bottom row: naive reconstructions using filtered back projection. The first two figures represent a frame-by-frame solution from measurements at frames 35 and 55. The third figure represents the static reconstruction from $f_{\text{bin}}$.
2.1 Motion Model
A motion model describes the relation between pixel intensities $u$ and the velocity flow $\mathbb{v}$ through an equation $r(u,\mathbb{v})=0$ in $\Omega_{T}$. Its choice is application-dependent, for instance, the continuity equation imposes a mass-preservation constraint, while the optical flow equation promotes no change in the intensities. These models are typically employed for the task of motion estimation, this is, determining the velocity flow $\mathbb{v}$ for the given image sequence $u$.
In this work, we focus on the well-known optical flow equation introduced in (3). It is derived from the brightness constancy assumption, which states that pixels keep constant intensity along their trajectory in time. This model poses a scalar equation for the $d$ components of the velocity field, leading to an underdetermined equation. This can be solved by considering a variational problem in $\mathbb{v}$ with a regularization term:
$\min_{\mathbb{v}}\mathcal{A}(r(u,\mathbb{v}))+\beta\mathcal{S}(\mathbb{v}),$ (4)
where $\mathcal{A}$ is a metric measuring how well the equation $r(u,\mathbb{v})=0$ is satisfied, $\mathcal{S}$ is a regularizer, and $\beta>0$ is the regularization parameter balancing both terms. This variational model was firstly introduced in [44] with $\mathcal{A}$ as the $L^{2}$-norm and $\mathcal{S}$ as the $L^{2}$-norm of the gradient. Since then, different norms and regularizers have been tried, for instance, in [45] the $L^{1}$-norm is used to impose the motion model, same as in [46] which employs the total variation for regularization.
2.2 Joint Image Reconstruction and Motion Estimation
To solve highly undersampled dynamic inverse problems, a joint variational problem is proposed in [8] where not only the dynamic process $u$ is sought, but also the underlying motion expressed in terms of a velocity field $\mathbb{v}$. The main hypothesis is that a joint reconstruction can enhance the discovery of both quantities, image sequence and motion, improving the final reconstruction compared to motionless models. Hence, the sought solution is a minimizer for the variational problem:
$\min_{u,\mathbb{v}}$ $\mathcal{D}(u,f_{t_{1}},\ldots,f_{t_{N_{T}}})+$ (5)
$\alpha\mathcal{R}(u)+\beta\mathcal{S}(\mathbb{v})+\gamma\mathcal{A}(r(u,\mathbb{v})),$
with $\alpha,\beta,\gamma>0$ being regularization parameters balancing the four terms. In [8], it is shown, among other things, how the pure motion estimation task of a noisy sequence can be enhanced by solving the joint task of image denoising and motion estimation.
This model was further employed for 2D+time problems in [10] and [11]. In the former, its application on dynamic CT is studied with sparse limited angles using both the $L^{1}$ and $L^{2}$-norms for the data fidelity term, with better results for the $L^{1}$-norm. In the latter, the same logic is used for dynamic cardiac MRI. In 3D+time domains, we mention [47] and [48] for dynamic CT and dynamic photoacoustic tomography respectively.
3 Methods
Depending on the nature of the noise, different data-fidelity terms can be considered. In this work, we consider Gaussian noise $\varepsilon$. To satisfy equation (2) at time $t$ we use an $L^{2}$ distance between predicted measurement and data $f$:
$\mathcal{D}_{t}(u,f):=\dfrac{1}{2}\|K_{t}[u_{t}]-f\|^{2}_{2}.$
The overall data-fidelity term in (5) is a mean over the measured times:
$\mathcal{D}(u,f_{t_{1}},\ldots,f_{t_{N_{T}}}):=\dfrac{1}{N_{T}}\displaystyle\sum_{i=1}^{N_{T}}\mathcal{D}_{t_{i}}(u,f_{t_{i}}).$ (6)
Since $u$ represents a natural image, a suitable choice for the regularizer $\mathcal{R}$ is the total variation in space to promote noiseless images and capture edges. For the motion model, we consider the optical flow equation (3), and to measure its discrepancy to 0 we use the $L^{1}$-norm; for the regularizer in $\mathbb{v}=(v_{1},\ldots,v_{d})^{T}$ we consider the total variation on each of its components:
$\mathcal{R}(u)$ $:=\int_{\Omega_{T}}\|\nabla u\|_{2},$ (7)
$\mathcal{A}(r(u,\mathbb{v}))$ $:=\int_{\Omega_{T}}|\partial_{t}u+\mathbb{v}\cdot\nabla u|,$
$\mathcal{S}(\mathbb{v})$ $:=\int_{\Omega_{T}}\sum_{j=1}^{d}\|\nabla v_{j}\|_{2}.$
For conciseness, we have omitted the dependency of the integrand on $(x,t)$.
We now describe how to solve the variational problem (5) numerically with neural fields. We proceed with a discretize-then-optimize approach. We also provide a brief description of the grid-based approach in [8] as we will compare it against our method.
3.1 Numerical evaluation with Neural Fields
We parametrize both the image and the motion with two independent neural fields $u_{\theta}$ and $\mathbb{v}_{\phi}$, respectively, with $\theta$ and $\phi$ denoting the weights and biases of the networks. Even though $u_{\theta}$ is a continuous and explicit representation, there is not a closed-form expression for its X-ray transform $K_{t}[(u_{\theta})_{t}]$. For this reason, the numerical evaluation of the forward operator is performed by first evaluating the network at points on a cartesian grid to get a grid-based representation at time $t$ as $(u_{\theta})_{t}:=\{u_{\theta}(x_{j},t)\}_{j=1,\ldots,N}$, and then applying our preferred implementation of the X-ray transform. In particular, we work with Tomosipo [49], a library that provides an integration of the ASTRA-toolbox [50, 51] with PyTorch, making it suitable for the optimization of neural networks. The data fidelity term in (6) requires first the evaluation of the network at $N\times N_{T}$ fixed grid-points to get the scene $\{u_{\theta}(x_{j},t_{i})\}_{i=1,\ldots,N_{T};j=1,\ldots,N}$, and, second, the application of the forward models $\{K_{t_{i}}\}_{i=1\ldots,N_{T}}$ to each frame. We call this the full-batch approach. This might be expensive and time-consuming. A common practice is to proceed with a mini-batch-like approach in time to speed up the optimization and avoid poor local minima. In this setting, at each iteration, we randomly sample $1\leq N_{b}\leq N_{T}$ frames, say, $\{i_{1},\ldots,i_{N_{b}}\}$, and the neural field is evaluated at the corresponding points to get the representation of the image at times $\{t_{i_{1}},\ldots,t_{i_{N_{b}}}\}$. Then, the forward model is applied on these frames only and the parameters are updated to minimize the difference between predicted data and the measured data $\{f_{t_{i_{1}}},\ldots,f_{t_{i_{N_{b}}}}\}$. This represents considerable benefits in terms of memory since the whole scene is never explicitly represented in the whole space-time grid and the cost of applying the forward operator on many frames is avoided, however, it adds variability during optimization.
Since neural fields are mesh-free, the regularization terms can be evaluated at any point of the domain. Additionally, derivatives can be computed through automatic differentiation. This motivates approximating the integrals for the regularizers in equation (7) via Monte Carlo by randomly sampling $N_{c}$ collocation points of the form $\{(x_{c},t_{c})\}_{c=1,\ldots,N_{c}}\subset\Omega_{T}$ using a Latin Hypercube Sampling strategy [52]. For ease of notation, we introduce the function $\eta$ defined as the integrand in the regularization terms:
$\eta(u,\mathbb{v},x,t):=$ $\alpha\|\nabla u(x,t)\|_{2}+\beta\sum_{j=1}^{d}\|\nabla v_{j}(x,t)\|_{2}$
$+\gamma|\partial_{t}u(x,t)+\mathbb{v}(x,t)\cdot\nabla u(x,t)|.$
At a given iteration, frames and collocation points are randomly sampled and the parameters $\theta$ and $\phi$ of the networks are updated by minimizing the following function as an unbiased estimator of the objective (5):
$\begin{array}[]{l}\dfrac{1}{N_{b}}\displaystyle\sum_{k=1}^{N_{b}}\mathcal{D}_{t_{i_{k}}}(u_{\theta},f_{t_{i_{k}}})+\dfrac{|\Omega_{T}|}{N_{c}}\displaystyle\sum_{c=1}^{N_{c}}\eta(u_{\theta},\mathbb{v}_{\phi},x_{c},t_{c}).\end{array}$
Finally, we recall that it is an open question how to choose $N_{c}$, the number of collocation points sampled at each iteration. One would like to sample as many points as possible to have a better approximation of the regularizer, however, this might be time-consuming and prohibitive in terms of memory because of the use of auto differentiation. Thus, we define the sampling rate ($SR$) as the ratio between $N_{c}$ and the amount of points $N_{T}\times N$ on the spatiotemporal grid:
$SR:=\dfrac{N_{c}}{N_{T}\times N}.$ (8)
3.2 Numerical evaluation with grid-based representation
In this section we briefly describe the numerical realization of the grid-based representation of (5) as in [8]. A uniform grid $\{(x_{j},t_{i})\}_{i=1,\ldots,N_{T};j=1,\ldots,N}\subset\Omega_{T}$ is assumed. Next, the quantities of interest are vectorized as $u\in\mathbb{R}^{N_{T}\times N}$, $\mathbb{v}\in\mathbb{R}^{N_{T}\times N\times d}$, such that, $u_{ij}$ denotes the value of $u$ at the point $(x_{j},t_{i})$. The evaluation of the data fidelity term is now straightforward using Tomosipo. For the regularization part finite difference schemes are employed to compute the corresponding derivatives. We let $D$ and $D_{t}$ denote the discretized gradients in space and time respectively (these could be forward or centred differences). Thus $(Du)\in\mathbb{R}^{N_{T}\times Nd}$, $(D_{t}u)\in\mathbb{R}^{N_{T}\times N}$, and $(Dv_{j})\in\mathbb{R}^{N_{T}\times N\times d}$ for $j=1,\ldots,d$. Thus, the regularizers in (7) are approximated as follows:
¯$\overline{\mathcal{R}}(u)$ $:=\dfrac{|\Omega_{T}|}{N_{T}N}\sum_{i=1}^{N_{T}}\|(Du)_{i}\|_{2,1}$
$\bar{\mathcal{A}}(r(u,v))$ $:=\dfrac{|\Omega_{T}|}{N_{T}N}\sum_{i=1}^{N_{T}}\|(D_{t}u)_{i}+\mathbb{v}_{i}\cdot(Du)_{i}\|_{1},$
$\overline{\mathcal{S}}(v)$ $:=\dfrac{|\Omega_{T}|}{N_{T}N}\sum_{i=1}^{N_{T}}\sum_{j=1}^{d}\|(Dv_{j})_{i}\|_{2,1},$
where $\|\cdot\|_{2,1}$ is a norm for vector fields, that first computes the 2-norm element-wise and then the 1-norm in space.
Using the previous, the variational problem (5) is discretized as
$\min_{u,v}$ $\mathcal{D}(u,f_{t_{1}},\ldots,f_{t_{N_{T}}})+$ (9)
$\alpha\overline{\mathcal{R}}(u)+\beta\overline{\mathcal{S}}(\mathbb{v})+\gamma\bar{\mathcal{A}}(r(u,\mathbb{v})),$
Due to the non-linearity present in the optical flow equation, this problem is not convex, however, it can be easily seen that it is biconvex, hence, in [8], the proposed optimization routine updates the current iteration $(u^{k},\mathbb{v}^{k})$ by alternating between the following two subproblems:
  • Problem in $u$. Fix $\mathbb{v}^{k}$ and update $u$ according to:
    $u^{k+1}=\operatorname*{arg\,min}_{u}$ $\mathcal{D}(u,f_{t_{1}},\ldots,f_{t_{N_{T}}})+$ (10)
    $\alpha\overline{\mathcal{R}}(u)+\gamma\bar{\mathcal{A}}(r(u,\mathbb{v}^{k}))$
  • Problem in $\mathbb{v}$. Fix $u^{k+1}$ and update $\mathbb{v}$ according to:
    $\mathbb{v}^{k+1}=\operatorname*{arg\,min}_{v}$ $\beta\overline{\mathcal{S}}(\mathbb{v})+\gamma\bar{\mathcal{A}}(r(u^{k+1},\mathbb{v})).$ (11)
Each subproblem is convex with non-smooth terms involved that can be solved using the Primal-Dual Hybrid Gradient (PDHG) algorithm [53]. We refer to [8] for further details.
4 Datasets
We study our method in a 2D+time setting. We employ two synthetic datasets, the two-square and the cardiac phantoms, and a real one, the STEMPO phantom [54]. For all the experiments the considered physical domain is the square $\Omega=[-1,1]^{2}$. The code will be made available on Github upon acceptance of the paper.
For the synthetic experiments, we define the ground-truth phantom $u:\Omega_{T}$ as follows:
  • Let $u_{0}:\Omega\to[0,1]$ be the initial frame, i.e., we let $u(\cdot,0):=u_{0}(\cdot)$.
  • Define $\varphi:\Omega_{T}\to\Omega$ describing the motion of the process. It takes a point $(x_{0},y_{0},t)\in\Omega_{T}$ and outputs $\varphi(x_{0},y_{0},t)\in\Omega$, the new position of $(x_{0},y_{0})$ at time $t$. For each time we can define the function $\varphi_{t}:\Omega\to\Omega$ by $\varphi_{t}(x_{0},y_{0})=\varphi(x_{0},y_{0},t)$. We require $\varphi_{t}$ to be a diffeomorphism for every $t\in[0,T]$. Hence we can define the trajectory of the point $(x_{0},y_{0})$ as $t\to\varphi_{t}(x_{0},y_{0})$.
  • Define $u(x,y,t):=u_{0}(\varphi_{t}^{-1}(x,y))$.
The phantom $u$ generated by the above procedure solves the optical flow equation with velocity field $\mathbb{v}=\frac{d}{dt}\varphi$. To mimic the continuous world, we define the ground truth at a high spatial resolution of $1024\times 1024$. Measurements are generated using Tomosipo assuming a camera with $M=64$ sensors. To avoid the inverse crime, during the reconstruction, we evaluate the neural field at a lower resolution spatial grid to get the discretized image as described in section 3.1.
For the synthetic datasets, we consider two sampling strategies, random and sequential. For the first, at each time a projection is taken along a random angle $\theta\in[0,2\pi)$; for the second, projections are acquired at fixed 9-degree intervals between consecutive frames. For STEMPO, both strategies are sequential, the first one with 32-degree intervals and the second with 4-degree intervals between frames.
4.1 Two-square phantom
Figure 3: Two-square phantom.
The first phantom is depicted in figure 3(a) with two squares moving within an ellipsis-shaped background. The inverse of the motion for the squares on the left and right are $\varphi_{1}^{-1}$ and $\varphi_{2}^{-1}$ respectively, each one given by the following expressions:
$\varphi_{1}^{-1}(x,y,t)$ $=\begin{pmatrix}x-\frac{t}{5}\cos(2\pi t)\\ y-\frac{3t}{4}\sin(2\pi t)\end{pmatrix},$
$\varphi_{2}^{-1}(x,y,t)$ $=\begin{pmatrix}x-0.3t\\ y-0.8t\end{pmatrix}.$
From this, the velocity fields are easily expressed as:
$v_{1}(x,y,t)$ $=\begin{pmatrix}\frac{1}{5}\cos(2\pi t)-\frac{2\pi t}{5}\sin(2\pi t)\\ \frac{3}{4}\sin(2\pi t)+\frac{3\pi t}{2}\cos(2\pi t)\end{pmatrix},$
$v_{2}(x,y,t)$ $=\begin{pmatrix}0.3\\ 0.8\end{pmatrix}.$
$v_{1}$ produces a spiral-like motion for the square on the left and $v_{2}$ a constant diagonal motion for the square on the right. These are depicted in the second row of figure 3(a) as follows: the colored boundary frame indicates the direction of the velocity field. The intensities of the image indicate the magnitude of the vector. As an example, the square on the right moves constantly up and slightly to the right during the motion.
For this phantom, we set the distance source-origin to 3, the distance source-detector to 5, and the detector size to 3.5 to specify the projection geometry. We consider $N_{T}=100$ frames, leading to a measurement array of dimension $64\times 100$. Measurements are further corrupted with Gaussian noise with standard deviation $\sigma=0.01$. See figure 3(b). During reconstruction, each frame is generated by evaluating the neural field at a spatial grid of resolution $64\times 64$.
Figure 4: Cardiac phantom.
4.2 Cardiac phantom
Our second phantom shown in figure 4(a) aims to mimic a heart-like motion. A slice of the heart is represented with an ellipse and three more circular-shaped structures are included. We also consider three cycles and set the final time to $T=3$. The motion is radial and described as follows:
$\varphi(x_{0},y_{0},t)=a(t)\begin{pmatrix}x_{0}\\ y_{0}\end{pmatrix},$
where the function $t\to a(t)$ is depicted in figure 4(c). As can be seen, it consists of three periods, with the first and third periods following the same pattern, while the second one depicts an intricate and irregular motion resembling an arrhythmia.
We set the same projection geometry as for the two-square phantom. For this phantom we consider $N_{T}=300$ frames, leading to a measurement array of dimension $64\times 300$. Measurements are further corrupted with Gaussian noise with standard deviation $\sigma=0.01$. See figure 4(b). During reconstruction, each frame is generated by evaluating the neural field at a spatial grid of resolution $64\times 64$.
4.3 STEMPO
In [54], it is introduced the Spatio-TEmporal Motor-Powered (STEMPO) dataset. It provides X-ray tomography data for a moving object. The dynamics of the object are controlled by a motor allowing for different sampling schemes to be performed. The phantom consists of two static objects and a square moving upwards from the bottom to the top at a constant speed.
The whole phantom is fully sampled at its initial and static state. A ground truth for the initial frame is then obtained by taking a filtered back projection and then setting to 0 those pixels below a given threshold. This is done to reduce the effect of the noise in the measurements and provide a clean background for the image. During the motion only one projection is acquired at a single time instance, collecting a total of 360 frames. Additionally, two sequential sampling schemes are provided, the first one takes projections every one degree, and the second one takes projections every 8 degrees. To generate the subsequent frames for the ground truth, the reconstructed initial frame is extrapolated according to the known motion given by the motor.
The acquired data is then binned at different factors. Here we consider a binning factor of 32, leading to each projection having 70 measurements. Given the simplicity of the motion, we add another layer of difficulty by uniformly downsampling the number of measured frames from 360 to 90. This leads to projections every 4 degrees and 32 degrees. For this phantom, we also consider a domain $[-1,1]^{2}\times[0,1]$, for which the real geometry is accordingly scaled, for instance, the distance source-origin is 9.88, the distance source-detector is 13.33, and the size of the detector is 2.69. During reconstruction, each frame is generated by evaluating the neural field at a spatial grid of resolution $70\times 70$.
Figure 5: STEMPO dataset.
5 Numerical Experiments
As mentioned before, in our numerical experiments, we use Tomosipo to compute the X-ray transform and its transpose. The architecture for the neural networks $u_{\theta}$ and $\mathbb{v}_{\phi}$ consists of an input layer of size 3, then Fourier feature mappings of the form $(x,t)\in\mathbb{R}^{3}\to(\Gamma_{1}(x),\Gamma_{2}(t))\in\mathbb{R}^{128}$, then three hidden layers of size 128 and we finish with an output layer of size 1 for $u_{\theta}$ and size $2$ for $\mathbb{v}_{\phi}$. We do not apply an activation function in the last layer. In particular, we do not impose the network $u_{\theta}$ to output non-negative values. The mappings are defined as $\Gamma_{1}(x):=(\sin(2\pi\mathbf{B_{x}}x),\cos(2\pi\mathbf{B_{x}}x))$ and $\Gamma_{2}(t):=(\sin(2\pi\mathbf{B_{t}}t),\cos(2\pi\mathbf{B_{t}}t))$, with the matrices $\mathbf{B_{x}}\in\mathbb{R}^{32\times 2}$ and $\mathbf{B_{t}}\in\mathbb{R}^{32\times 1}$ having non-trainable entries $(\mathbf{B_{x}})_{ij}\sim\mathcal{N}(0,\sigma_{x}^{2})$ and $(\mathbf{B_{t}})_{ij}\sim\mathcal{N}(0,\sigma_{t}^{2})$. The sinusoidal functions act element-wise. We found $\sigma_{x}=\sigma_{t}=0.1$ to give the best results for the two-square and the STEMPO datasets, and $\sigma_{x}=0.1,\sigma_{t}=0.5$ for the cardiac dataset. We try two batch settings, the full-batch $N_{b}=N_{T}$ and the mini-batch $N_{b}=1$. In the full-batch setting we optimize for 150,000 iterations, while in the mini-batch setting, we optimize for 10,000 epochs (leading to $10,000\times N_{T}$ iterations). We set the sampling rate (8) at $SR=0.1$, leading to 40,960, 122,880, and 44,100 collocation points being randomly sampled in each iteration for the two-square, cardiac, and STEMPO datasets, respectively. We use the Adam optimizer with a learning rate of $10^{-3}$ in all cases. For the grid-based approach, each subproblem, (10) and (11), are run for 2,000 iterations and this alternation is repeated 5 times. The PSNR values reported in this section are computed using the resolution during reconstruction. Since the ground truth is defined at a higher resolution in space, we downsample it using pooling average. Experiments are performed on a Nvidia Tesla V100 GPU with 16GB [55].
5.1 Unregularized versus regularized neural fields
Figure 6: Ablation study on $\gamma$ for neural fields. Evolution of PSNR every 100 iterations during optimization with batch size $N_{b}=N_{T}$.
Figure 7: Evolution of PSNR every 100 iterations for the unregularized ($\gamma=0$) and the proposed regularized solution $(\gamma=10^{-2})$ with neural fields.
Figure 8: Varying batch size $N_{b}\in\{1,100\}$ for $\gamma=0$ and $\gamma=10^{-2}$. Evolution of PSNR against time in hours during optimization of the neural field.
Figure 9: Reconstruction (left) with neural fields and error (right) of the two-square phantom for the random sampling strategy at frames 35, 55, and 75 (out of 100).
In this section, we study the effect of optimizing the neural field with an explicit motion regularizer and compare it against the purely implicitly regularized solution. For this purpose, we set the regularization parameters $\alpha=\beta=0$ and perform an ablation study on $\gamma\in\{10^{-4},10^{-3},10^{-2},10^{-1},1\}$ in the three datasets with random sampling and measure its performance in terms of PSNR.
In figure 8 we show the evolution of PSNR during optimization for the three datasets. This figure highlights the role played by $\gamma$ in the reconstruction. On the one hand, for the lowest value, $\gamma=10^{-4}$, it can be seen a semi-convergence behavior where approximately after 40,000 epochs the neural field starts fitting the noise in the measurements and the reconstruction quality decreases steadily throughout the optimization. On the other hand, for the highest value, $\gamma=1$, the regularization term is too strongly imposed and a completely static image is obtained. For the two-square phantom, $\gamma=10^{-2}$ gives the best result in terms of PSNR achieving a value of 34.41 after 150,000 iterations; for the cardiac phantom, $\gamma=10^{-3}$ performs the best during most of the optimization, however, its quality starts decreasing after 60,000 iterations and the case $\gamma=10^{-2}$ achieves a higher PSNR of 28.03 at the last iteration; for STEMPO, we can see that both $\gamma=10^{-2}$ and $\gamma=10^{-1}$ achieve a similar PSNR but the former evolves more quickly and attains a value of 23.51. We therefore propose a regularized solution using $\gamma=10^{-2}$ for the three datasets.
We continue our study by comparing the unregularized case $\gamma=0$ against the proposed regularized one. This is shown in figure 8. In all cases, $\gamma=0$ shows degradation on the reconstruction early during optimization, after around 30,000 iterations. Its regularized counterpart $\gamma=10^{-2}$ is superior across almost all the iterations. More importantly, this regularizer stabilizes the reconstruction and prevents the neural field from fitting the noise in the measurements.
We also compare the full-batch setting $N_{b}=N_{T}$ against the commonly used mini-batch setting with $N_{b}=1$, and for $\gamma\in\{0,10^{-2}\}$. We recall that for $N_{b}=1$ optimization is done for 10,000 epochs and results are displayed every 100 epochs. Results for the evolution of PSNR versus time are shown in figure 8. A batch size of 1 speeds up the optimization achieving a higher PSNR in less time. It suffers however from an unstable behaviour with the PSNR varying for almost 2 points every 100 epochs. This means we can stop at a poor reconstruction without a reliable stopping criteria.
We show the obtained reconstructions at their best PSNR and the corresponding error against the ground truth at different frames for the two-square, cardiac, and STEMPO datasets in figures 9, 11, and 11, respectively. The top rows show the full-batch regularized case $N_{b}=N_{T},\gamma=10^{-2}$; the middle rows show the mini-batch regularized case $N_{b}=1,\gamma=10^{-2}$; the bottom rows show the full-batch unregularized case $N_{b}=N_{T},\gamma=0$. A relevant advantage in the mini-batch setting is that it is more likely to capture edges in less time since more iterations can be taken, while its full-batch counterpart $N_{b}=N_{T}$ shows blurry edges. We also summarize the PSNR values obtained for the different models in table 1 highlighting the role of $\gamma$.
Parameters Datasets
$\gamma$ $N_{b}$ Two-square Cardiac STEMPO
$10^{-2}$ $N_{T}$ 34.52 28.09 23.91
$10^{-2}$ 1 34.28 25.32 23.91
0 $N_{T}$ 25.58 23.03 22.36
0 1 27.46 21.88 23.16
Table 1: Maximum PSNR attained during optimization obtained by neural fields with $\alpha=\beta=0$ and varying the motion regularization parameter $\gamma$ and the batch size $N_{b}$.
Figure 10: Reconstruction (left) with neural fields and error (right) of the cardiac phantom for the random sampling strategy at frames 1, 122, and 250 (out of 300).
Figure 11: Reconstruction (left) with neural fields and error (right) of the STEMPO phantom for the sequential 32-degree sampling strategy at frames 1, 40, and 60 (out of 90).
5.2 Regularized neural fields versus grid-based method.
Figure 12: Neural field versus grid-based for the two-square phantom.
Figure 13: Neural field versus grid-based for the cardiac phantom.
Two-square
$(\alpha,\beta,\gamma)=(10^{-3},10^{-4},10^{-3})$
Cardiac
$(\alpha,\beta,\gamma)=(10^{-4},10^{-4},5\times 10^{-3})$
Grid-based Neural Field Grid-based Neural Field
PSNR 27.09 32.92 17.55 29.77
Data-fidelity $6.25\times 10^{-5}$ $7.41\times 10^{-5}$ $1.22\times 10^{-4}$ $8.51\times 10^{-5}$
$\alpha\mathcal{R}(u)$ $5.1\times 10^{-4}$ $5.04\times 10^{-4}$ $1.81\times 10^{-4}$ $1.25\times 10^{-4}$
$\beta\mathcal{S}(v)$ $4.6\times 10^{-5}$ $1.05\times 10^{-6}$ $4.14\times 10^{-4}$ $4.3\times 10^{-5}$
$\gamma\mathcal{A}(r(u,v))$ $1.3\times 10^{-4}$ $2.23\times 10^{-4}$ $1.12\times 10^{-4}$ $5.21\times 10^{-3}$
Final loss $7.53\times 10^{-4}$ $8.03\times 10^{-4}$ $8.29\times 10^{-4}$ $5.46\times 10^{-3}$
Table 2: Comparing the grid-based method against neural fields for the two-square and cardiac phantoms. We show PSNR and the terms from the loss function. In particular, the regularization terms for the neural field are obtained by evaluating it at the whole spatiotemporal grid.
Figure 14: Comparing grid-based method against neural fields for the two-square phantom with sequential sampling. Reconstruction (left) and error (right).
Figure 15: Comparing grid-based method against neural fields for the STEMPO phantom with sequential 4-degree sampling. Reconstruction (left) and error (right).
We now compare explicitly regularized neural fields against the grid-based method outlined in section 3.2. For this approach, it was found that the choice $(\alpha,\beta,\gamma)=(10^{-3},10^{-4},10^{-3})$ led to the best results in terms of PSNR, achieving a value of 27.09. We solve the same variational problem with neural fields by choosing the same regularization parameters, in which case the PSNR achieved is 32.92. Figure 12(a) shows reconstructions at different frames. To appreciate the time evolution we also show a $x-t$ slice view in figure 12(b) at the horizontal center $y=0$, where it becomes clear that the grid-based solution struggles at capturing regularity in time.
The same experiment is performed for the cardiac dataset, for which the best reconstruction was attained at $(\alpha,\beta,\gamma)=(10^{-4},10^{-4},5\times 10^{-3})$ achieving a PSNR of 17.55 for the grid-based method and 29.77 for the neural field (notice that this value is higher than the PSNR attained for $\gamma=10^{-2}$ in the previous section). In this case, the grid-based method completely fails at the reconstruction task and the rapid motion of this phantom cannot be captured. Figure 13(b) shows a $y-t$ slice at the vertical center $x=0$ and it can be seen that the inner circle at the bottom barely moves. The neural field on the other hand is still able to show regularity in time. This shows that neural fields can outperform the grid-based solution even for the choice of regularization parameters that led to the best behavior for the grid-based method.
We report the PSNR and loss values for both experiments in table 2. For the neural field, the regularizers are estimated by evaluating it on the same cartesian grid as for the grid-based representation. There we can see that the grid-based achieves a lower loss for both experiments, in particular, it attains a lower value for the data fidelity and the optical flow terms. We highlight that a very low optical flow is related to static motion as observed in the cardiac phantom in figure 13(b).
We finish our study by comparing both methods at the more challenging problem of sequential sampling. We try the sequential 9-degree sampling for the two-square phantom and the 4-degree sampling for STEMPO. For the two-square phantom, we use the regularization parameters that gave the best results in the previous section, namely, $(\alpha,\beta,\gamma)=(10^{-3},10^{-4},10^{-3})$ for the grid-based method achieving a PSNR of 22.65. For the neural field, we use $(\alpha,\beta,\gamma)=(0,0,10^{-2})$ and the method achieves a PSNR of 26.42. Results are shown in figure 15. There it can be seen that both methods perform worse than in the random sampling, but still, the neural field reconstruction is better than the grid-based one which shows large errors, for instance, in frame 60. For STEMPO we found $(\alpha,\beta,\gamma)=(10^{-4},10^{-4},10^{-1})$ to give the best reconstruction for the grid-based method with a PSNR of 14.24; for the neural field, we set $(\alpha,\beta,\gamma)=(0,0,10^{-2})$, leading to a PSNR of 15.18. Results can be seen in figure 15. In this case, both methods can capture the static part of the image but fail at the reconstruction of the moving square due to the sampling scheme.
The bad reconstructions shown in the first two rows of figure 15 motivated us to try different strategies for the neural field. In particular, we observe a different behavior for the choice $(\alpha,\beta,\gamma)=(10^{-3},10^{-4},10^{-3})$: during optimization, the neural field fits the static part but after 15,000 iterations approximately, the optical flow error begins to increase, promoting a better reconstruction of the dynamic part but, at the same time, worsening the static part. This behavior can be captured during the optimization since we have access to the optical flow loss. Thus, we try an adaptive routine, where the value of $\gamma$ is increased if the optical flow error increases as well. We increase the value of $\gamma$ from $10^{-3}$ to $10^{-2}$ when the optical flow error increases. This reconstruction is shown in the third row of figure 15. The PSNR achieved is now 17.39 and both static and dynamic parts are better captured. This shows that neural fields can potentially solve the challenging sequential sampling case with more dedicated optimization routines. We leave this for future research. We do not try such a strategy for the grid-based case, hence, we cannot conclude that this is a particular benefit of the neural field representation.
6 Conclusion
This work considers neural fields for dynamic CT image reconstruction from finely resolved in time but severely undersampled in space measurements as commonly encountered in applications, e.g. cardiac CT. Neural fields are particularly suitable for this task: their continuity allows for a coherent representation in both time and space; their resolution independence enables us to query it at multiple frames, and, together with their differentiability, facilitate the numerical evaluation of the PDE-based motion regularizer. We studied how to enhance the neural field reconstruction by making use of the optical flow equation: constraining the neural field to this physically feasible motion meant a significant improvement with respect to the state-of-the-art both in terms of neural fields as well as traditional grid-based representations.
[…]
Acknowledgments PA is supported by a scholarship from the EPSRC Centre for Doctoral Training in Statistical Applied Mathematics at Bath (SAMBa), under the project EP/S022945/1. MJE acknowledges support from the EPSRC (EP/S026045/1, EP/T026693/1, EP/V026259/1, EP/Y037286/1) and the European Union Horizon 2020 research and innovation programme under the Marie Skodowska-Curie grant agreement REMODEL. The authors gratefully acknowledge the University of Bath’s Research Computing Group (doi.org/10.15125/b6cd-s854) for their support in this work.